Worked with Jack to isolate, by resetting power supply and cycling power but could not consistently reset supply.
Called Mark for direction and he recommended changing the local control board. We did this and multiple resets showed this to clear problem.
Reset power supply and left Hall Took beam for 5 minutes and power supply faulted with earth leakage fault.
Re-entered the Hall and tried boosting lead flow as per Heidi’ recommendation to Jack Via phone.
According to Heidi earth leakage fault= lead flow fault
Boosted lead flow to 90 slpm
We also noted that after the earth leakage fault the power supply would arbitrarily ramp to ~800 amps the settle to requested current. Mark says this may be a problem in the mother board.
Left Hall and set up strip charts to track amperage and lead flows
Heidi came in and pointed out that the lead flows had dropped at about 5:00 and that that corilated with the first power supply trip. She also pointed out that the target J T valve had been opened at about the same time.
According to Heidi the process meter for the lead flow may be failing due to radiation damage and adding to the lead flow problem.
Called cryo on call to troubleshoot lead flow
After having a stable magnet with the boosted for 40 min. went back to hall noted all lead flows and warm return pressure. Normal pressure is 1.5 psi pressure found was 2 psi
Received call from Jonathon Creel that the return pressure was indeed linked to the JT valve change and that the recovery compressors where at full capacity and he recommended we turn back the JT valve a bit to stabilize the system.
Called Bryan and put him in touch with Jonathan to coordinate JT valve.
Hopefully this will get us through the weekend.
We should put in a request to software to change this error message
A copy of this log entry has been emailed to: fansler@jlab.org